OpenStudy (anonymous):

You drop a stone down a well and hear the echo 8.9 s later. if it take 09 s for the echo to travel up the well, approximately how deep is the well?

OpenStudy (anonymous):

y=1/2*gt^2-----and----y=1/2(9.8)(9*9)------>y=397m